Heat can be removed from water at 0C and atmospheric pressure without causing the water to freeze, if done with little disturbance of the water. Suppose the water is cooled to –5C before ice began to form. What is the change in entropy per unit mass occurring during the sudden freezing that then takes place?
1
Expert's answer
2019-12-18T11:28:28-0500
"\Delta S =C_p \ln{\frac{T_2}{T_1}}"

"\Delta S =4200 \ln{\frac{273-5}{273}}=-77.6\frac{J}{kgK}"